OpenStudy (anonymous):

One person can do a certain job in fifteen minutes, and another person can do the same job in thirty minutes. How many minutes will it take them to do the job together? 5 min 10 min 20 min

OpenStudy (anonymous):

first person does work in 1 minutes = 1/15 of total work second person does work in 1 minutes = 1/30 of total work both will do in one minute = 1/15 + 1/30 = 1/10 of total work. so in ten minutes they will complete the work.

OpenStudy (anonymous):

Thank you very much!